OpenStudy (kj4uts):

Use the discriminant to determine the number of real roots of... x^2-5x-7=0 x^2-2x+1=0 Choices are: 0 1 2

Nnesha (nnesha):

what is discriminant ? do you know ?

OpenStudy (kj4uts):

Well my work is (I just want to check): x^2-5x-7=0 -5^2-4(1)(-7)=3>0 2 roots x^2-2x+1=0 -2^2-4(1)(1)=-8<0 1 root

OpenStudy (kj4uts):

b^2-4ac

OpenStudy (kj4uts):

x^2-5x-7=0 2 roots x^2-2x+1=0 1 root

Nnesha (nnesha):

\(\huge\color{reD}{\rm b^2-4ac}\) `Discriminant` you can use this to find if the equation is factorable or not if ` b^2-4ac > 0` then there are 2 real zeros if ` b^2-4ac = 0` then there is one real root if ` b^2-4ac < 0` then you will get two complex roots (no -x-intercept)

Nnesha (nnesha):

you should use discriminant

Nnesha (nnesha):

ohh wait nmv

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @KJ4UTS Well my work is (I just want to check): x^2-5x-7=0 -5^2-4(1)(-7)=3>0 2 roots x^2-2x+1=0 -2^2-4(1)(1)=-8<0 1 root \(\color{blue}{\text{End of Quote}}\) b^2 so it wohuld be (-2)^2 -4(1)(1)

Nnesha (nnesha):

so (-2)^2 -4 = ?

OpenStudy (kj4uts):

-8

Nnesha (nnesha):

no (-2)^2 is same as -2 times -2

Nnesha (nnesha):

or in other words when you take `even` power of negative exponent you will get positive answer always !

OpenStudy (kj4uts):

0

Nnesha (nnesha):

no \[\color{ReD}{(-2)^2}-4(1)(1)\] i'm just talkign about (-2)^2

Nnesha (nnesha):

talking *

Nnesha (nnesha):

yes right

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \(\huge\color{reD}{\rm b^2-4ac}\) `Discriminant` you can use this to find if the equation is factorable or not if ` b^2-4ac > 0` then there are 2 real zeros if ` b^2-4ac = 0` then there is one real root if ` b^2-4ac < 0` then you will get two complex roots (no -x-intercept) \(\color{blue}{\text{End of Quote}}\) now read this when b^2-4ac =0 how many roots will u get ?

OpenStudy (kj4uts):

x^2-2x+1=0 -2^2-4(1)(1)=0=0 Which would still be 1 root?

Nnesha (nnesha):

yes right

Nnesha (nnesha):

now what about first one ? b^2-4ac = ??

OpenStudy (kj4uts):

Was my work right -5^2-4(1)(-7)=3>0 2 roots

Nnesha (nnesha):

no that's not correct ....

OpenStudy (kj4uts):

53

OpenStudy (kj4uts):

I think what I was doing wrong for both was not putting parenthesis (-5)^2 for both problems into my calculator but 53>0 therefore I think it is still 2 roots?

Nnesha (nnesha):

yes right

OpenStudy (kj4uts):

Ok thank you @Nnesha for your time and help, and for also pointing out where I went wrong :)

Nnesha (nnesha):

np and yes you should put the parentheses (-2)^2 like this :=) good job!